Carlos works 8.25 h and receives $102.30 pay. What will he receive at the same rate if he works if he works 14 h?
A) $189.10 B) $204.60 C) $173.60 D) $158.10
I thought I would get the answer by doing this... $102.30 divided by $8.25= 12.4 h so $8.25 multiplied by 14 h = $115.50, but as you can see this answer is not one of my options. What am I doing wrong?

Answer by stanbon(75887)

Carlos works 8.25 h and receives $102.30 pay. What will he receive at the same rate if he works if he works 14 h?
A) $189.10 B) $204.60 C) $173.60 D) $158.10
---------------------------------------
EQUATION:
x/14 = 102.3/8.25
x = 14*102.3/8.25
x = $173.60
=================
Cheers,
Stan H.